How long will it take for $2800 to be worth $7000 at a rate of 7.8%/ yearly compounded annually
1
Expert's answer
2020-06-15T18:12:41-0400

F = P( 1 + r/n)nt

7000 = 2800(1 + 0.078/1)1*t

7000/2800= (1 + 0.078)t

log(2.5) = t log (1.078)

log(2.5)/ log (1.078) = t

t = 12.20yrs